# GCoT: Chain-of-Thought Prompt Learning for Graphs

### v1

@@ 5.1. Experimental Setup
Datasets. We conduct experiments on eight widely used benchmark datasets, spanning citation networks, e-commerce, protein structures, and molecular graphs. Cora (McCallum et al., 2000), Citeseer (Sen et al., 2008), and Pubmed (Sen et al., 2008) are citation networks, each consisting of a single graph. In these datasets, nodes represent academic papers and edges denote citation relationships between them. Photo (Shchur et al., 2018) is an e-commerce co-purchase network derived from Amazon’s photography-related product category, where nodes correspond to products and edges indicate frequently co-purchased items. PROTEINS (Borgwardt et al., 2005) is a dataset of protein structures. In each graph, nodes correspond to secondary structures, and edges capture spatial or sequential relationships within the amino acid sequence. MUTAG (Rossi and Ahmed, 2015), BZR (Rossi and Ahmed, 2015), and COX2 (Rossi and Ahmed, 2015) are molecular graph datasets, representing nitroaromatic compounds, ligands associated with benzodiazepine receptors, and COX2 contains molecular structures related to cyclooxygenase-2 inhibitors, respectively. A detailed summary of these datasets is presented in Table 1, with further descriptions in Appendix A.
@@
Baselines. We compare GCoT with state-of-the-art methods across three categories: (1) Supervised GNNs: GCN (Kipf and Welling, 2017) and GAT (Veličković et al., 2018) are trained directly on downstream labels in a fully supervised manner, without any pre-training. (2) Graph pre-training models: DGI/InfoGraph [^1] DGI is originally designed for node-level tasks, while InfoGraph extends it to graph-level classification. In our experiments, we apply DGI to node classification and InfoGraph to graph classification. (Veličković et al., 2018; Sun et al., 2020) and GraphCL (You et al., 2020) adopt a “pre-train, fine-tune” strategy. These models first perform self-supervised pre-training using unlabeled graphs and are later fine-tuned for downstream tasks, where a classifier is trained with few-shot labels while keeping the pre-trained encoder frozen in single step. (3) Graph prompt learning models: ProG (Sun et al., 2023), GPF (Fang et al., 2024), GPF+ (Fang et al., 2024), and GraphPrompt (Liu et al., 2023) employ self-supervised pre-training followed by prompt tuning. Unlike the fine-tuning methods, these methods leverage a unified task template, and train task-specific prompts in single step for downstream adaptation. Further descriptions of these baselines and implementation details are shown in Appendix B &  C, respectively.
@@
Downstream tasks and evaluation. We perform experiments on two downstream tasks: node classification and graph classification. Both tasks follow an $m$-shot classification setup, where for each class, we randomly select $m$ instances (nodes or graphs) as labeled examples. In our main results, we set $m=1$ for both node and graph classification tasks. Additionally, to examine the robustness of our method, we vary $m$ within the range $1\leq m\leq 10$, allowing us to analyze performance under different few-shot scenarios. We construct 100 independent $m$-shot tasks for each classification setting through repeated sampling. Each task is evaluated using five different random seeds, resulting in a total of 500 experimental runs per task type. We report both the mean and standard deviation across these results. We evaluate performance using accuracy, in line with previous studies (Wang et al., 2020; Liu et al., 2021a, 2023).
@@ 5.2. Performance Evaluation
We first evaluate one-shot classification tasks. Then, we examine the effect of increasing the number of shots on model performance.
One-shot performance. We present the results for one-shot node and graph classification tasks in Tables 2. We observe that: (1) GCoT consistently outperforms most baseline methods, surpassing the best competitor by margins of up to 11.05% in node classification and 6.03% in graph classification. These results underscore the advantage of incorporating multiple reasoning steps to better adapt to diverse tasks. (2) Standard graph prompt learning approaches—ProG, GPF, GPF+, and GraphPrompt—exhibit significantly weaker performance compared to GCoT. Their limitations arise from relying on a single-step inference process that lacks iterative refinement of the final answer. This contrast highlights the effectiveness of our chain-of-thought prompting, which enables step-by-step inference and captures individualized learning patterns for each node.
Few-shot performance. To evaluate the impact of labeled data availability on GCoT’s performance, we vary the number of shots in both node and graph classification tasks. The results, shown in Fig. 3, compare GCoT against several competitive baselines. First, GCoT consistently outperforms the baselines, particularly in low-shot scenarios (e.g., $m\leq 5$), where labeled data is scarce. Second, as the number of labeled samples increases (e.g., $m>5$), all methods generally exhibit improved performance, which aligns with expectations. Nevertheless, GCoT remains highly competitive, often achieving the best results. Note that on certain datasets, such as PROTEINS, performance across methods tends to show high variance. A possible reason is that this dataset exhibits greater variability in graph sizes compared to other datasets: The standard deviation of graph sizes in PROTEINS is 45.78, whereas other datasets fall within the range of 4.04 to 15.29. This may contribute to unstable performance. Despite this, GCoT demonstrates greater robustness than the competing methods. For the rest of the experiments, we focus on the one-shot classification setting.
Figure 3. Impacts of shots on node and graph classification.
@@ 5.3. Ablation Study and Visualization
GCoT-L1 61.45±15.82 59.94±14.84 56.54±14.12 54.71±8.57
GCoT-L2 66.95±16.07 63.05±15.13 57.68±13.84 54.77±8.81
GCoT-L3 64.39±15.76 69.27±18.20 57.85±16.10 56.22±8.45
GCoT 66.09±15.17 70.57±17.10 58.75±15.42 56.24±8.60
Figure 4. Visualization of output embedding space of nodes. Different colors represents different classes.
@@ 5.4. Heterophily Sensitivity
To examine the robustness of GCoT on heterophilic graphs (Pei et al., 2020; Yu et al., 2025), we conduct one-shot node classification on heterophilic datasets (Wisconsin (Pei et al., 2020) and Squirrel (Rozemberczki et al., 2021)). We provide the detailed description of these datasets in Appendix D. As shown in Table 4, GCoT consistently outperforms other competing baselines, further validating its effectiveness. These results indicates that the iterative inference process in GCoT successfully generalizes across both homophilic and heterophilic graphs.
Table 4. Accuracy (%) evaluation on heterophilic datasets.
Methods Wisconsin Squirrel
DGI 28.04 $\pm$ 6.47 20.00 $\pm$ 1.86
GraphCL 29.85 $\pm$ 8.46 21.42 $\pm$ 2.22
ProG 32.95 $\pm$ 11.42 20.61 $\pm$ 3.01
GPF 30.29 $\pm$ 10.04 19.89 $\pm$ 3.19
GPF+ 29.26 $\pm$ 9.61 20.93 $\pm$ 3.98
GraphPrompt 29.96 $\pm$ 9.40 21.22 $\pm$ 1.80
GCoT 33.04 $\pm$ 9.71 22.03 $\pm$ 4.94
@@ 5.5. Flexibility of Graph Prompting Methods
To evaluate the flexibility and robustness of GCoT, we evaluate its performance using various standard graph prompting methods as the task prompt. Specifically, we integrate ProG (Sun et al., 2023), GPF (Fang et al., 2024), GPF+ (Fang et al., 2024), and GraphPrompt (Liu et al., 2023) into our framework. The results for both node and graph classification on four datasets are presented in Table 5. Across nearly all cases, GCoT consistently outperforms its single-step prompting counterparts, regardless of the task prompt employed. These findings highlight the robustness of our approach and further confirm the advantage of chain-of-thought prompting over previous single-step prompting.
Table 5. Evaluation of GCoT with different standard prompting methods.
Prompting Thinking Node classification Graph classification
@@
ProG - 51.35±13.92 64.07±14.74 51.99±14.50 52.73±6.57
GCoT 59.62±14.20 72.19±14.92 58.42±15.39 55.82±8.79
GPF - 57.35±13.87 57.61±11.26 56.55±13.95 53.05±7.62
GCoT 60.28±15.69 66.24±12.61 57.66±13.11 56.98±8.30
GPF+ - 64.91±15.49 57.85±10.86 56.81±12.93 54.58±8.70
GCoT 62.71±15.18 65.97±11.09 58.54±13.29 55.45±9.20
GraphPrompt - 54.56±12.56 64.85±10.72 55.44±12.56 53.80±7.93
GCoT 55.83±13.69 72.73±16.18 58.28±15.47 55.32±8.43
“-” refers to standard graph prompting without our CoT design.
5.6. Hyperparameter Analysis for Condition-Net
In our experiments, we implement the condition-net as a two-layer MLP with a bottleneck design. To examine its effect on performance, we vary the hidden dimension $s$ and present the results in Fig. 5. Our findings reveal that $s=32$ generally yields competitive results for node classification, and $s=8$ for graph classification, leading us to adopt these settings in our experiments. A smaller $s$ may restrict the model’s representational capacity, limiting its effectiveness, whereas a larger $s$ introduces additional trainable parameters and increases the likelihood of overfitting in few-shot learning scenarios.
@@ 5.6. Hyperparameter Analysis for Condition-Net
5.7. Effect of Varying Number of Inference Steps
We further vary the number $K$ of inference steps to analyze their impact, with the results presented in Fig. 6. We observe that for both node and graph classification tasks, $K=2$ generally yields optimal performance. Hence, we adopt $K=2$ in our experiments. For Proteins, $K=3$ achieves the best result, likely due to its inherent complexity as described in Sect. 5.2.
@@ 6. Conclusions
In this paper, we propose GCoT, the first CoT prompting framework for graphs. We define an inference step with three substages: prompt-based inference, thought construction, and thought conditioned prompt learning. Specifically, we first feed the prompt modified query into the pre-trained encoder, and then construct a thought by fusing the hidden embeddings from each layer of the pre-trained graph encoder. To guide the subsequent inference step, we generate a series of prompts conditioned on the thought from the previous step. By repeating the above inference steps, GCoT obtain the answer. Finally, we conduct extensive experiments on eight public datasets, demonstrating that GCoT significantly outperforms a range of state-of-the-art baselines.
@@ A. Further Descriptions of Datasets
In this section, we provide more comprehensive descriptions of the benchmark datasets used in our experiments, as summarized in Table 1, in the following.
@@
  • Cora [^2] https://github.com/shchur/gnn-benchmark/raw/master/data/npz/cora.npz (McCallum et al., 2000) is a citation network composed of 2,708 research papers in the field of computing, each classified into one of seven categories. The network consists of 5,429 citation links between papers. Each paper is represented by a binary word vector, where each entry indicates the presence or absence of a word from a predefined vocabulary of 1,433 unique terms.
@@
  • Citeseer [^3] https://github.com/shchur/gnn-benchmark/raw/master/data/npz/citeseer.npz (Sen et al., 2008) includes 3,312 computer science publications, categorized into six distinct classes, separate from those in Cora. The citation network contains 4,732 edges. Each document is encoded as a binary word vector that captures the presence or absence of words from a dictionary comprising 3,703 unique terms.
@@
  • PubMed [^4] https://github.com/shchur/gnn-benchmark/raw/master/data/npz/pubmed.npz (Sen et al., 2008) is a citation network of 19,717 biomedical articles related to diabetes, divided into three categories. The network includes 44,338 citation edges. Unlike Cora and Citeseer, each document is represented by a TF/IDF-weighted word vector derived from a dictionary of 500 unique terms.
@@
  • Photo [^5] https://github.com/shchur/gnn-benchmark/raw/master/data/npz/amazon_electronics_photo.npz (Shchur et al., 2018) consists of 7,487 photography-related products, each assigned to one of eight categories. The co-purchase network contains 119,043 edges, where connections indicate products frequently bought together. Each product is described by a feature vector extracted from metadata and customer reviews, with category labels corresponding to product types.
@@
  • MUTAG [^6] https://huggingface.co/datasets/graphs-datasets/MUTAG (Debnath et al., 1991) is a dataset of nitroaromatic compounds aimed at predicting their mutagenic effects on Salmonella typhimurium. Each compound is modeled as a graph, where nodes represent atoms with categorical labels (encoded as one-hot vectors) based on atom types, and edges depict the chemical bonds connecting them. The dataset comprises 188 molecular graphs with 7 unique node types.
@@
  • BZR [^7] https://chrsmrrs.github.io/datasets/docs/datasets/ (Rossi and Ahmed, 2015) consists of 405 molecular graphs representing ligands that interact with benzodiazepine receptors. Each molecule is treated as an independent graph and is classified into one of two categories.
@@
  • COX2888https://chrsmrrs.github.io/datasets/docs/datasets/ (Rossi and Ahmed, 2015) includes 467 molecular structures of cyclooxygenase-2 inhibitors. In this dataset, nodes correspond to atoms, while edges define chemical bonds—which may be single, double, triple, or aromatic. The molecules are divided into two distinct classes.
@@
  • PROTEINS [^9] https://chrsmrrs.github.io/datasets/docs/datasets/ (Borgwardt et al., 2005) is a dataset of protein structure graphs that encode both biochemical and structural properties. In this dataset, nodes represent secondary structural elements, while edges capture connectivity based on spatial proximity or amino acid sequence adjacency. Each node falls into one of three categories, and graphs are classified into two broader groups.

### v2

@@ 5.1. Experimental Setup
Datasets. We conduct experiments on eight widely used benchmark datasets, spanning citation networks, e-commerce, protein structures, and molecular graphs. Cora (McCallum et al., 2000), Citeseer (Sen et al., 2008), and Pubmed (Sen et al., 2008) are citation networks, each consisting of a single graph. In these datasets, nodes represent academic papers and edges denote citation relationships between them. Photo (Shchur et al., 2018) is an e-commerce co-purchase network derived from Amazon’s photography-related product categories, where nodes correspond to products and edges indicate frequently co-purchased items. PROTEINS (Borgwardt et al., 2005) is a dataset of protein structures. In each graph, nodes correspond to secondary structures, and edges capture spatial or sequential relationships within the amino acid sequence. MUTAG (Rossi and Ahmed, 2015), BZR (Rossi and Ahmed, 2015), and COX2 (Rossi and Ahmed, 2015) are molecular graph datasets, representing nitroaromatic compounds, ligands associated with benzodiazepine receptors, and molecular structures related to cyclooxygenase-2 inhibitors, respectively. These datasets are summarized in Table 1, with further descriptions in Appendix A.
@@
Baselines. We compare GCoT with state-of-the-art methods across three categories. (1) Supervised GNNs: GCN (Kipf and Welling, 2017) and GAT (Veličković et al., 2018) are trained directly on downstream labels in a supervised manner, without any pre-training. (2) Graph pre-training models: DGI/InfoGraph [^2] DGI is originally designed for node-level tasks, while InfoGraph extends it to graph-level classification. In our experiments, we apply DGI to node classification and InfoGraph to graph classification. (Veličković et al., 2018; Sun et al., 2020) and GraphCL (You et al., 2020) adopt a “pre-train, fine-tune” strategy. They first perform self-supervised pre-training using unlabeled graphs and are later fine-tuned for downstream tasks, where a classifier is trained with few-shot labels while keeping the pre-trained encoder frozen. (3) Graph prompt learning models: ProG (Sun et al., 2023), GPF (Fang et al., 2024), GPF+ (Fang et al., 2024), and GraphPrompt (Liu et al., 2023) employ self-supervised pre-training followed by prompt tuning. Unlike the fine-tuning methods, these methods leverage a unified task template, and train task-specific prompts for downstream adaptation. Further descriptions of these baselines are presented in Appendix B.
@@
Downstream tasks and evaluation. We perform experiments on two downstream tasks: node classification and graph classification. Both types of task follow an $m$-shot classification setup, where for each class, we randomly select $m$ instances (nodes or graphs) as labeled examples. The remaining instances are treated as the test set. In our main results, we set $m=1$ for both node and graph classification tasks. Additionally, to examine the robustness of our method, we vary $m$ within the range $[1,10]$, allowing us to analyze the performance under different few-shot scenarios. We construct 100 independent $m$-shot tasks for each type of classification through repeated sampling. Each task is evaluated using five different random seeds, resulting in a total of 500 experimental runs per task type. We adopt accuracy as the performance metric and report both the mean and standard deviation across these runs, in line with previous studies (Wang et al., 2020; Liu et al., 2021a, 2023).
5.2. Implementation Details
We outline key settings for the baselines and GCoT.
Baseline settings. We utilize the official codes for all open-source baselines. Each model was tuned based on the settings recommended in their respective work to achieve optimal performance.
For the baseline GCN (Kipf and Welling, 2017), we employ a 2-layer architecture, and set the hidden dimensions to 64. For GAT (Veličković et al., 2018), we employ a 2-layer architecture and set the hidden dimension to 64. Additionally, we apply 8 attention heads in the first GAT layer. For DGI (Veličković et al., 2018), we utilize a 1-layer GCN as the backbone and set the hidden dimensions to 256. Additionally, we employ PReLU as the activation function. For InfoGraph (Sun et al., 2020), a 1-layer GCN is used as the backbone, with its hidden dimensions set to 256. For GraphCL (You et al., 2020), a 1-layer GCN is also employed as its backbone, with the hidden dimensions set to 256. Specifically, we select edge dropping as the augmentations, with a default augmentation ratio of 0.2. For GraphPrompt (Liu et al., 2023), a 3-layer GCN is used as the backbone for all datasets, with the hidden dimensions set to 256. For GPF and GPF+ (Fang et al., 2024), we employ a 3-layer GCN as the backbone for all datasets. The hidden dimensions are set to 256. For ProG (Sun et al., 2023), we employ a 2-layer GCN as the backbone for all datasets. The hidden dimensions are set to 100.
GCoT settings. For our proposed GCoT, we utilize a 3-layer GCN as the backbone for all datasets, with the hidden dimensions set to 256. We set the number of inference step as 2 for node classification task, and 3 for graph classification task. The condition-net is implemented as a two-layer MLP with a bottleneck architecture (Yu et al., 2025c, b). Its input dimension is 256, while the hidden dimension is set to 32 for node classification and 8 for graph classification.
Environment. All experiments were conducted on Ubuntu 22.04.2, using a machine equipped with AMD EPYC 7742 64-core processors and NVIDIA GeForce RTX 3090 (24 GB) GPUs.
@@ 5.3. Performance Evaluation
We first evaluate one-shot classification tasks. Then, we examine the model performance as the number of shots increases. Note that for the other experiments in Sect. 5.4 and thereafter, we adopt the one-shot setting.
One-shot performance. We present the results for one-shot node and graph classification in Table 2. We make several major observations, as follows.
(1) GCoT consistently outperforms the baseline methods across both node and graph classification, demonstrating its overall advantage and robustness.
(2) Supervised methods tend to underperform compared to others, as they do not leverage any pre-trained model. On the other hand, graph pre-training models achieve improved performance through self-supervised pre-training on unlabeled graphs.
(3) Standard graph prompt learning approaches—ProG, GPF, GPF+, and GraphPrompt—often outperform fine-tuning of the pre-trained models, due to their alignment between pre-training and downstream objectives, as well as their parameter efficiency. However, they still fall short compared to GCoT, due to their single-step inference process that lacks iterative refinement of the final answer. This highlights the effectiveness of our CoT-style prompting, which enables step-by-step inference.
Few-shot performance. To investigate the effect of labeled data size on performance, we vary the number of shots in both node and graph classification tasks. The results are presented in Fig. 3, comparing GCoT against several competitive baselines. We make the following observations.
(1) GCoT generally outperforms the baselines on both node and graph classification tasks, particularly in low-shot settings (e.g., $m\leq 5$), where labeled data are limited.
(2) As the number of shots increases (e.g., $m>5$), all methods generally exhibit improved performance, as expected. Nevertheless, GCoT remains highly competitive, often achieving the best or near-best results.
(3) On the PROTEINS dataset, the performance of all methods tends to fluctuate more. A possible reason is that this dataset exhibits greater variability in graph sizes compared to other datasets: The standard deviation of graph sizes in PROTEINS is 45.78, whereas other datasets fall within the range of 4.04 to 15.29. This may contribute to the unstable performance. Despite this, GCoT demonstrates greater robustness than the competing methods.
Figure 3. Impact of labeled data size (number of shots) on node and graph classification.
@@ 5.4. Ablation Study and Visualization
To thoroughly evaluate the contribution of CoT-style prompting to graph models, we perform an ablation study and visualization.
[…]
@@
GCoT-L1 57.18±14.34 63.31±10.05 56.54±14.12 54.71±8.57
GCoT-L2 57.00±14.48 63.20±10.08 57.68±13.84 54.77±8.81
GCoT-L3 57.01±14.66 63.33±10.05 57.85±16.10 56.22±8.45
GCoT 59.67±15.51 64.43± 9.96 58.75±15.42 56.24±8.60
Figure 4. Visualization of node and thought embeddings. Different colors represent different node classes.
@@ 5.5. Heterophily Sensitivity
To examine the robustness of GCoT on heterophilic graphs (Yu et al., 2025c), we conduct one-shot node classification on Wisconsin (Pei et al., 2020) and Squirrel (Rozemberczki et al., 2021). Specifically, Wisconsin is a network of 251 nodes representing webpages, with 199 edges indicating hyperlink connections among them. Node features are constructed using a bag-of-words approach based on webpage content. Squirrel (Rozemberczki et al., 2021) consists of 5,201 Wikipedia pages discussing predefined topics. Nodes represent individual webpages, while edges capture 217,073 hyperlink connections between them. Node features are constructed using key informative nouns extracted from the text content of the Wikipedia pages. As shown in Table 4, GCoT outperforms other competitive baselines on heterophilic node classification. These results indicate that our CoT-style prompting can generalize across both homophilic and heterophilic graphs.
Table 4. Accuracy (%) evaluation of node classification on heterophilic graphs.
Methods Wisconsin Chameleon
DGI 28.04 $\pm$ 06.47 19.33 $\pm$ 4.57
GraphCL 29.85 $\pm$ 08.46 27.16 $\pm$ 4.31
ProG 28.99 $\pm$ 11.14 25.90 $\pm$ 3.95
GPF 32.18 $\pm$ 11.36 25.98 $\pm$ 3.71
GPF+ 32.02 $\pm$ 11.17 26.12 $\pm$ 3.75
GraphPrompt 31.48 $\pm$ 05.18 25.36 $\pm$ 3.99
GCoT 32.80 $\pm$ 11.59 27.96 $\pm$ 3.89
@@ 5.6. Flexibility of Graph Prompting Methods
To evaluate the flexibility and robustness of GCoT, we evaluate its performance using various standard graph prompting methods as the standard prompt. We integrate ProG (Sun et al., 2023), GPF (Fang et al., 2024), GPF+ (Fang et al., 2024), and GraphPrompt (Liu et al., 2023) into our framework. Specifically, ProG, GPF, and GPF+ are employed to modify the input features at the first inference step, whereas GraphPrompt is utilized to modify the output embeddings at the final step. The results for both node and graph classification on four datasets are presented in Table 5. Across all cases, GCoT outperforms its standard prompting counterparts, highlighting the robustness and flexibility of our CoT design.
Table 5. Accuracy (%) evaluation of GCoT with different standard prompting methods.
Prompting Node classification Graph classification
@@
ProG 50.59±14.64 63.07±11.96 51.99±14.50 52.73±6.57
with GCoT 52.62±15.27 64.01±11.27 58.42±15.39 55.82±8.79
GPF 57.60±13.88 55.63±10.96 56.55±13.95 53.05±7.62
with GCoT 58.98±15.28 62.63±10.01 57.66±13.11 56.98±8.30
GPF+ 57.42±13.87 57.16±10.99 56.81±12.93 54.58±8.70
with GCoT 59.35±15.37 62.18±10.49 58.54±13.29 55.45±9.20
GraphPrompt 54.25± 9.38 63.11±10.01 55.44±12.56 53.80±7.93
with GCoT 58.40±14.74 63.79± 9.40 58.28±15.47 55.32±8.43
In each group, the first row represents a standard prompting method, and the second row represents the standard prompting coupled with our CoT design.
5.7. Hyperparameter Analysis
Next, we study the sensitivity of two key hyperparameters, namely, the hidden dimension of the condition-net, and the number of inference steps performed.
Hidden dimension of condition-net. In our experiments, we implement the condition-net as a two-layer MLP with a bottleneck design. To examine the effect of the MLP architecture, we vary its hidden dimension $s$ and present the results in Fig. 5. The results reveal that $s=32$ generally yields strong results for node classification, and $s=8$ for graph classification. A smaller $s$ may restrict the model’s representational capacity, limiting its effectiveness, whereas a larger $s$ introduces additional trainable parameters and increases the risk of overfitting in few-shot scenarios.
@@ 5.7. Hyperparameter Analysis
Number of inference steps. We further investigate the influence of the number of inference steps, $K$, with the results summarized in Fig. 6. For node classification, we observe that $K=2$ generally yields the best performance, whereas $K=3$ is optimal for graph classification. This discrepancy may stem from the fact that graph classification relies more on global structural information, which may be better captured with more inference steps. Notably, for the Proteins dataset, $K=7$ achieves the highest performance, likely due to its intrinsic complexity and variability, as discussed in Sect. 5.3.
@@
5.8. Computational Efficiency Analysis
To evaluate the efficiency of GCoT, we perform one-shot node classification on a large-scale dataset called CLUSTER (Dwivedi et al., 2023), which comprises 1,172,320 nodes and 43,038,630 edges. The accuracy evaluation is shown in Table 6, whereas inference time comparisons with strong baselines are reported in Table 7.
Table 6. Accuracy (%) evaluation of node classification on the large-scale CLUSTER dataset.
Methods DGI GraphCL GraphPrompt GPF+ GCoT
Accuracy 21.27 23.82 22.18 24.27 27.54
While GCoT maintains higher accuracy on the large-scale dataset, we observe additional inference time due to its multi-step inference on the test set. Such a limitation of CoT has also been observed in NLP. In our case, inference time increases roughly linearly with the number of steps, which is acceptable considering the significant accuracy gains.
Table 7. Inference time (milliseconds) on the test set of a node classification task.
Methods Cora Pubmed MUTAG CLUSTER
GraphPrompt 3.90 5.62 5.85 955.25
GPF+ 5.25 5.92 4.79 713.08
GCoT 7.95 13.95 11.44 2340.00
@@ 6. Conclusions
In this paper, we proposed GCoT, the first CoT prompting framework for text-free graphs. Specifically, we introduced the notion of step-by-step inference, where each inference step consists of three substages: prompt-based inference, thought construction, and thought-conditioned prompt learning. Concretely, we first feed the prompt-modified query graph into a pre-trained graph encoder, and then construct a thought by fusing the hidden embeddings from all layers of the pre-trained encoder to capture hierarchical structural knowledge. Subsequently, conditioned on the thought, we generate a series of node-specific prompts to guide the next inference step. After multiple such inference steps, GCoT makes a prediction on the final “answer.” Lastly, we conducted extensive experiments on eight public datasets, demonstrating that GCoT significantly outperforms a range of state-of-the-art baselines.
Acknowledgments
This research / project is supported by the Ministry of Education, Singapore, under its Academic Research Fund Tier 2 (Proposal ID: T2EP20122-0041). Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of the Ministry of Education, Singapore.
@@ A. Further Descriptions of Datasets
  • Cora (McCallum et al., 2000) is a citation network composed of 2,708 research papers in the field of computing, each classified into one of seven categories. The network consists of 5,429 citation links between papers. Each paper is represented by a binary word vector, where each entry indicates the presence or absence of a word from a predefined vocabulary of 1,433 unique terms.
@@
  • Citeseer (Sen et al., 2008) includes 3,312 computer science publications, categorized into six distinct classes, separate from those in Cora. The citation network contains 4,732 edges. Each document is encoded as a binary word vector that captures the presence or absence of words from a dictionary comprising 3,703 unique terms.
@@
  • PubMed (Sen et al., 2008) is a citation network of 19,717 biomedical articles related to diabetes, divided into three categories. The network includes 44,338 citation edges. Unlike Cora and Citeseer, each document is represented by a TF/IDF-weighted word vector derived from a dictionary of 500 unique terms.
@@
  • Photo (Shchur et al., 2018) consists of 7,487 photography-related products, each assigned to one of eight categories. The co-purchase network contains 119,043 edges, where connections indicate products frequently bought together. Each product is described by a feature vector extracted from metadata and customer reviews, with category labels corresponding to product types.
@@
  • MUTAG (Debnath et al., 1991) is a dataset of nitroaromatic compounds aimed at predicting their mutagenic effects on Salmonella typhimurium. Each compound is modeled as a graph, where nodes represent atoms with categorical labels (encoded as one-hot vectors) based on atom types, and edges depict the chemical bonds connecting them. The dataset comprises 188 molecular graphs with 7 unique node types.
@@
  • BZR (Rossi and Ahmed, 2015) consists of 405 molecular graphs representing ligands that interact with benzodiazepine receptors. Each molecule is treated as an independent graph and is classified into one of two categories.
@@
  • COX2 (Rossi and Ahmed, 2015) includes 467 molecular structures of cyclooxygenase-2 inhibitors. In this dataset, nodes correspond to atoms, while edges define chemical bonds—which may be single, double, triple, or aromatic. The molecules are divided into two distinct classes.
@@
  • PROTEINS (Borgwardt et al., 2005) is a dataset of protein structure graphs that encode both biochemical and structural properties. In this dataset, nodes represent secondary structural elements, while edges capture connectivity based on spatial proximity or amino acid sequence adjacency. Each node falls into one of three categories, and graphs are classified into two broader groups.